whats 8/15 x 9/4 x 3/8

Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

216/480=9/20

Answer by Alan3354(69443)

whats 8/15 x 9/4 x 3/8
--------
It's 8*9*3/15*4*8 = 216/480
= 9/20